# EvolMPNN: Predicting Mutational Effect on Homologous Proteins by Evolution Encoding

### v1

@@ 4.4 Final Embedding and Optimisation
After obtaining protein $\mathcal{P}_{i}$’s residue embeddings $\mathbf{R}_{i}$ and sequence embedding $\mathbf{Z}^{\textsc{P}}_{i}$, we summarise its residue embeddings as a vector $\mathbf{Z}^{\textsc{R}}_{i}=\textsc{MEAN Pool}(\mathbf{R}_{i})$. The final protein embedding summarises the protein sequence information and evolution information as the comprehensive embedding $\mathbf{Z}_{i}=\textsc{Concat}(\mathbf{Z}^{\textsc{P}}_{i},\mathbf{Z}^{\textsc{R}}_{i})$ and the final prediction is computed as $\hat{\mathbf{Y}}_{i}=\mathbf{Z}_{i}\mathbf{W}^{\textsc{Final}}$ where $\mathbf{W}^{\textsc{Final}}\in\mathbb{R}^{d\times\theta}$, $\theta$ is the number of properties to predict. Afterwards, we adopt a simple and common strategy, similar to [5], to solve the protein property prediction tasks. Specifically, we adopt the MSELoss ($\mathcal{L}$) to measure the correctness of model predictions on training samples against ground truth labels. The objective of learning the target task is to optimise model parameters to minimise the loss $\mathcal{L}$ on this task. The framework of EvolMPNN is summarised in Algorithm 1 in Appendix A.
4.5 Extensions on Observed Graph
EvolGNN. We compute the evolution information as Eq. 3. The evolution information can be easily integrated into the pipeline of message-passing neural networks, as an additional structural coefficient [51]:
$\mathbf{m}^{(\ell)}_{a}$ $=\textsc{Aggregate}^{\mathcal{N}}(\{\mathbf{A}_{ij},\underbrace{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}\mathbf{d}_{ij}}}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}\text{Evol. info.}}},\mathbf{H}^{(\ell-1)}_{j}\,|\,j\in\mathcal{N}(i)\}),$ (7)
$\mathbf{m}^{(\ell)}_{i}$ $=\textsc{Aggregate}^{\mathcal{I}}(\{\mathbf{A}_{ij},\underbrace{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}\mathbf{d}_{ij}}}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}\text{Evol. info.}}}\,|\,j\in\mathcal{N}(i)\})\,\mathbf{H}^{(\ell-1)}_{i},$
$\mathbf{Att}^{h}(\mathbf{H}^{(\ell)})=\textsc{Softmax}(\frac{\mathbf{H}^{(\ell)}\mathbf{W}^{\ell,h}_{Q}(\mathbf{H}^{(\ell)}\mathbf{W}^{\ell,h}_{K})^{\textsc{T}}}{\sqrt{d}}+\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\textsc{MEAN Pool}(\{\mathbf{R}_{i}\}_{i=1,2,\dots,M})}}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\text{Evol. info.}}}),$ (8)
4.6 Theoretical Analysis
Anchor Selection. Inspired by [52], we adopt Bourgain’s Theorem [53] to guide the random anchor number ($k$) of the evolution encoding layer. Briefly, support by a constructive proof (Theorem 2 [54]) of Bourgain Theorem (Theorem 1), only $k=O(\log^{2}M)$ anchors are needed to ensure the resulting embeddings are guaranteed to have low distortion (Definition 1), in a given metric space $(\mathcal{M},\mathrm{F}_{\textsc{Dist}})$. EvolMPNN can be viewed as a generalisation of the embedding method of Theorem 2, where $\mathrm{F}_{\textsc{Dist}}(\cdot)$ is generalised via message passing functions (Eq 3-Eq. 6). Therefore, Theorem 2 offers a theoretical guide that $O(\log^{2}M)$ anchors are needed to guarantee low distortion embedding. Following this principle, EvolMPNN choose $k=\log^{2}M$ random anchors, denoted as $\{S_{j}\}_{j=1,2,\dots,\log^{2}M}$, and we sample each protein in $\mathcal{M}$ independently with probability $\frac{1}{2^{j}}$. Detailed discussion and proof refer to Appendix B.
5 Experimental Study
In this section, we empirically study the performance of EvolMPNN. We validate our model on three benchmark homologous protein family datasets and evaluate the methods on nine data splits to consider comprehensive practical use cases. Our experiments comprise a comprehensive set of state-of-the-art methods from different categories. We additionally demonstrate the effectiveness of two extensions of our model, EvolGNN and EvolFormer, with different input features. We conclude our analysis studying the influence of some hyper-parameters and investigating the performance of EvolMPNN on high mutational mutants.
Table 1: Datasets splits, and corresponding statistics; if the split comes from a benchmark paper, we report the corresponding citation.
Landscape Split # Total #Train #Valid #Test
AAV [14] 2-vs-Rest [4] 82,583 28,626 3,181 50,776
7-vs-Rest [4] 82,583 63,001 7,001 12,581
Low-vs-High [4] 82,583 42,791 4,755 35,037
GB1 [16] 2-vs-Rest [4] 8,733 381 43 8,309
3-vs-Rest [4] 8,733 2,691 299 5,743
Low-vs-High [4] 8,733 4,580 509 3,644
Fluorescence [15] 2-vs-Rest 54,025 12,712 1,413 39,900
3-vs-Rest [5] 54,025 21,446 5,362 27,217
Low-vs-High 54,025 44,082 4,899 5,044
Datasets and Splits. We perform experiments on benchmark datasets of several important protein engineering tasks, including AAV, GB1 and Fluorescence, and generate three splits on each dataset. Data statistics are summarised in Table 1. The split $\lambda$-vs-Rest indicates that we train models on wild-type protein and mutants of no more than $\lambda$ mutations, while the rest are assigned to test. The split Low-vs-High indicates that we train models on sequences with target value scores equal to or below wild-type, while the rest are assigned to test. For more details refer to Appendix C.
Implementation. We follow the PEER benchmark settings [^4] https://github.com/DeepGraphLearning/PEER_Benchmark, including train and test pipeline, model optimisation and evaluation method (evaluation is Spearman’s $\rho$ metric), adopted in [5] to make sure the comparison fairness. For the baselines, including feature engineer, protein sequence encoder and pre-trained PLM, we adopt the implementation provided by benchmark Torchdrug [41] and the configurations reported in [5]. For the GNN-based baselines, which require predefined graph structure and protein features, we construct $K$-NN graphs [60], with $K=\{5,10,15\}$, and report the best performance. As features, we use the trained sequence encoder, which achieves better performance, used also in our method. In addition, we adopt ESM-1b as the residue encoder on GB1 dataset and adopt One-Hot encoding on AAV and Fluorescence datasets to speed up the training process. All experiments are conducted on two NVIDIA GeForce RTX 3090 GPUs with 24 GB memory, and we report the mean performance of three runs with different random seeds. We present more details in Appendix D. Note that we do not report results that take more than 48 hours due to our limited computation resources.
5.1 Effectiveness
Table 2: Quality in terms Spearman’s $\rho$ correlation with target value. NA indicates a non-applicable setting. * Used as a feature extractor with pre-trained weights frozen. $\dagger$ Results reported in [4, 5]. - Can not complete the training process within 48 hours on our devices. Top-2 performances of each split are marked as bold and underline.
Category Model Dataset
AAV GB1 Fluorescence
BLOSUM62 NA NA NA 0.128 0.005 -0.127 NA NA NA
DDE ${}^{\dagger}$ 0.636 0.158 ${}^{\dagger}$ 0.816 0.306 0.690 ${}^{\dagger}$ 0.159
Moran ${}^{\dagger}$ 0.398 0.069 ${}^{\dagger}$ 0.589 0.193 0.445 ${}^{\dagger}$ 0.046
Protein Seq. Encoder LSTM ${}^{\dagger}$ 0.608 0.308 ${}^{\dagger}$0.002 -0.007 0.256 ${}^{\dagger}$ 0.207
Transformer ${}^{\dagger}$ 0.748 0.304 ${}^{\dagger}$ 0.877 0.474 0.250 ${}^{\dagger}$ 0.161
CNN ${}^{\dagger}$ 0.730 0.406 ${}^{\dagger}$ 0.857 0.515 0.805 ${}^{\dagger}$ 0.249
ResNet ${}^{\dagger}$ 0.733 0.223 ${}^{\dagger}$ 0.542 0.396 0.594 ${}^{\dagger}$ 0.243
Pre-trained PLM ProtBert ${}^{\dagger}$ ${}^{\dagger}$ 0.866 0.308 0.451 ${}^{\dagger}$
ProtBert* ${}^{\dagger}$ 0.507 0.277 ${}^{\dagger}$ 0.619 0.164 0.403 ${}^{\dagger}$ 0.161
ESM-1b ${}^{\dagger}$ ${}^{\dagger}$ 0.878 0.386 0.804 ${}^{\dagger}$
ESM-1b* ${}^{\dagger}$ 0.573 0.241 ${}^{\dagger}$ 0.605 0.178 0.528 ${}^{\dagger}$ 0.091
ESM-1v* 0.533 0.580 0.171 0.359 0.632 0.180 0.562 0.563 0.070
ESM-2* 0.475 0.581 0.199 0.422 0.632 0.189 0.501 0.511 0.084
GAT 0.821 0.741 0.369 0.757 0.873 0.508 0.768 0.667 0.208
GraphTransf. 0.827 0.749 0.389 0.753 0.876 0.548 0.780 0.678 0.231
NodeFormer 0.827 0.741 0.393 0.757 0.877 0.543 0.794 0.677 0.213
Ours EvolMPNN 0.835 0.757 0.433 0.768 0.881 0.584 0.809 0.684 0.228
EvolMPNN outperforms all baselines on 8 of 9 splits. Table 2 summarises performance comparison on AAV, GB1 and Fluorescence datasets. EvolMPNN achieves new state-of-the-art performance on most splits of three datasets, with up to $6.7\%$ improvements to baseline methods. This result vindicates the effectiveness of our proposed design to capture evolution information for homologous protein property prediction. Notably, GNN-based methods that utilise manually constructed graph structure do not enter top-2 on 9 splits and two Transformer structure models, i.e., GraphTransformer and NodeFormer, often outperform such methods. It can be understood since homology graph construction is a challenging biomedical task [26], the simple $K$-NN graph construction is not an effective solution.
Large-scale PLM models are dominated by simple models. Surprisingly, we find that smaller models, such as CNN and ResNet, can outperfoerm large ESM variaants pre-trained on million- and billion-scale sequences. For instance, ESM-1v has about 650 million parameters and is pre-trained on around 138 million UniRef90 sequences [25]. Yet, CNN outperforms ESM-1v on three splits of Fluorescence dataset. This indicates the necessity of designs targeting specifically the crucial homologous protein engineering task.
Model Split
2-vs-R. 3-vs-R. L.-vs-H.
Best Baseline 0.757 0.878 0.548
EvolMPNN ($\Phi^{\textsc{OH}}$) 0.766 0.877 0.553
EvolGNN ($\Phi^{\textsc{OH}}$) 0.764 0.866 0.536
EvolFormer ($\Phi^{\textsc{OH}}$) 0.764 0.868 0.537
EvolMPNN ($\Phi^{\textsc{PLM}}$) 0.768 0.881 0.584
EvolGNN ($\Phi^{\textsc{PLM}}$) 0.767 0.879 0.581
EvolFormer ($\Phi^{\textsc{PLM}}$) 0.766 0.879 0.575
5.2 Analysis of Performance
Figure 4: EvolMPNN performance on AAV’s 2-vs-Rest (a) and Low-vs-High (b) splits, with different hyper-parameters.
Influence of hyper-parameter settings on EvolMPNN. We present in Figure 4 a group of experiments to study the influence of some hyper-parameters on EvolMPNN, including the number of attention heads, embedding dimension and the number of layers of residue encoder and evolution encoder. EvolMPNN exhibits stable performance on different hyper-parameter settings.
6 Conclusion and Future Work
We propose Evolution-aware Message Passing Neural Network (EvolMPNN), that integrates both protein sequence information and evolution information by means of residues to predict the mutational effect on homologous proteins. Empirical and theoretical studies show that EvolMPNN and its extended variants (EvolGNN and EvolFormer) achieve outstanding performance on several benchmark datasets while retaining reasonable computation complexity. In future work, we intend to incorporate 3D protein structure information towards general-purpose homologous protein models.
Acknowledgments
Appendix A Algorithm
@@ 24 $\mathbf{Z}^{\textsc{R}}_{i}=\textsc{MEAN Pool}(\mathbf{R}_{i})$
We summarise the process of Evolution-aware Message Passing Neural Network (EvolMPNN) in Algorithm 1. Given a protein set =$\mathcal{M}=\{\mathcal{P}_{i}\}_{i=1,2,\dots,M}$ and each protein sequence $\mathcal{P}_{i}$ contains a residue set $\{r_{j}\}_{j=1,2,\dots,N}$. For each protein $\mathcal{P}_{i}$, we first initialise protein sequence ($\mathbf{H}_{i}$) and residue embeddings ($\mathbf{X}_{i}$) (line 1-2). After, the residue embeddings are empowered with positional encoding ($\Phi_{\textsc{Pos}}$) to get $\hat{\mathbf{X}}_{i}$ (line 3). Such a design will help us to record the position of a mutation occurring in the protein sequence in the following steps. Then, we update residue embedding based on a contact map, which records the chemical reactions between residues after folding into a 3D structure (line 4-10). Furthermore, we aggregate evolution-aware embeddings by means of updated residue embeddings (line 11- line 17) and integrate them with protein sequence embeddings to empower them with evolution-aware semantics (line 18-21). Finally, we merge protein sequence and residue embeddings as the final protein embeddings, which contain comprehensive information and make predictions based on them (line 22- line 23).
Appendix B Theoretical Analysis

### v2

@@ 4.4 Final Embedding and Optimisation
After obtaining protein $\mathcal{P}_{i}$’s residue embeddings $\mathbf{R}_{i}$ and sequence embedding $\mathbf{Z}^{\textsc{P}}_{i}$, we summarise its residue embeddings as a vector $\mathbf{Z}^{\textsc{R}}_{i}=\textsc{MEAN Pool}(\mathbf{R}_{i})$. The final protein embedding summarises the protein sequence information and evolution information as the comprehensive embedding $\mathbf{Z}_{i}=\textsc{Concat}(\mathbf{Z}^{\textsc{P}}_{i},\mathbf{Z}^{\textsc{R}}_{i})$ and the final prediction is computed as $\hat{\mathbf{Y}}_{i}=\mathbf{Z}_{i}\mathbf{W}^{\textsc{Final}}$ where $\mathbf{W}^{\textsc{Final}}\in\mathbb{R}^{d\times\theta}$, $\theta$ is the number of properties to predict. Afterwards, we adopt a simple and common strategy, similar to [44], to solve the protein property prediction tasks. Specifically, we adopt the MSELoss ($\mathcal{L}$) to measure the correctness of model predictions on training samples against ground truth labels. The objective of learning the target task is to optimise model parameters to minimise the loss $\mathcal{L}$ on this task. The framework of EvolMPNN is summarised in Algorithm 1.
@@ 24 $\mathbf{Z}^{\textsc{R}}_{i}=\textsc{MEAN Pool}(\mathbf{R}_{i})$
4.5 Extensions on Observed Graph
EvolGNN. We compute the evolution information as Eq. 3. The evolution information can be easily integrated into the pipeline of message-passing neural networks, as an additional structural coefficient:
$\mathbf{m}^{(\ell)}_{a}$ $=\textsc{Aggregate}^{\mathcal{N}}(\{\mathbf{A}_{ij},\underbrace{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{d}_{ij}}}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\text{Evol. info.}}},\mathbf{H}^{(\ell-1)}_{j}\,|\,j\in\mathcal{N}(i)\}),$ (7)
$\mathbf{m}^{(\ell)}_{i}$ $=\textsc{Aggregate}^{\mathcal{I}}(\{\mathbf{A}_{ij},\underbrace{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{d}_{ij}}}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\text{Evol. info.}}}\,|\,j\in\mathcal{N}(i)\})\,\mathbf{H}^{(\ell-1)}_{i},$
$\mathbf{Att}^{h}(\mathbf{H}^{(\ell)})=\textsc{Softmax}(\frac{\mathbf{H}^{(\ell)}\mathbf{W}^{\ell,h}_{Q}(\mathbf{H}^{(\ell)}\mathbf{W}^{\ell,h}_{K})^{\textsc{T}}}{\sqrt{d}}+\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textsc{MEAN Pool}(\{\mathbf{R}_{i}\}_{i=1,2,\dots,M})}}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{Evol. info.}}}),$ (8)
4.6 Theoretical Analysis
Anchor Selection. Inspired by [45], we adopt Bourgain’s Theorem [3] to guide the random anchor number ($k$) of the evolution encoding layer. Briefly, support by a constructive proof (Theorem 0.A.2 [24]) of Bourgain Theorem (Theorem 0.A.1), only $k=O(\log^{2}M)$ anchors are needed to ensure the resulting embeddings are guaranteed to have low distortion (Definition 1), in a given metric space $(\mathcal{M},\mathrm{F}_{\textsc{Dist}})$. EvolMPNN can be viewed as a generalisation of the embedding method of Theorem 0.A.2, where $\mathrm{F}_{\textsc{Dist}}(\cdot)$ is generalised via message passing functions (Eq 3-Eq. 6). Therefore, Theorem 0.A.2 offers a theoretical guide that $O(\log^{2}M)$ anchors are needed to guarantee low distortion embedding. Following this principle, EvolMPNN choose $k=\log^{2}M$ random anchors, denoted as $\{S_{j}\}_{j=1,2,\dots,\log^{2}M}$, and we sample each protein in $\mathcal{M}$ independently with probability $\frac{1}{2^{j}}$. Detailed discussion and proof refer to Appendix 0.A.
5 Experimental Study
In this section, we empirically study the performance of EvolMPNN. We validate our model on three benchmark homologous protein family datasets and evaluate the methods on nine data splits to consider comprehensive practical use cases. Our experiments comprise a comprehensive set of state-of-the-art methods from different categories. We additionally demonstrate the effectiveness of two extensions of our model, EvolGNN and EvolFormer, with different input features. We conclude our analysis by studying the influence of some hyper-parameters and investigating the performance of EvolMPNN on high mutational mutants.
Table 1: Dataset splits and corresponding statistics; if the split comes from a benchmark paper, we report the corresponding citation.
Landscape Split # Total #Train #Valid #Test
AAV [5] 2-vs-Rest [8] 82,583 28,626 3,181 50,776
7-vs-Rest [8] 82,583 63,001 7,001 12,581
Low-vs-High [8] 82,583 42,791 4,755 35,037
GB1 [42] 2-vs-Rest [8] 8,733 381 43 8,309
3-vs-Rest [8] 8,733 2,691 299 5,743
Low-vs-High [8] 8,733 4,580 509 3,644
Fluorescence [36] 2-vs-Rest 54,025 12,712 1,413 39,900
3-vs-Rest [44] 54,025 21,446 5,362 27,217
Low-vs-High 54,025 44,082 4,899 5,044
Datasets and Splits. We perform experiments on benchmark datasets of several important protein engineering tasks, including AAV, GB1 and Fluorescence, and generate three splits on each dataset. Data statistics are summarised in Table 1. The split $\lambda$-vs-Rest indicates that we train models on wild-type protein and mutants of no more than $\lambda$ mutations, while the rest are assigned to test. The split Low-vs-High indicates that we train models on sequences with target value scores equal to or below wild-type, while the rest are assigned to the test.
Implementation. We follow the PEER benchmark settings [^4] https://github.com/DeepGraphLearning/PEER_Benchmark, including train and test pipeline, model optimisation and evaluation method (evaluation is Spearman’s $\rho$ metric), adopted in [44] to make sure the comparison fairness. For the baselines, including feature engineer, protein sequence encoder and pre-trained PLM, we adopt the implementation provided by benchmark Torchdrug [48] and the configurations reported in [44]. For the GNN-based baselines, which require predefined graph structure and protein features, we construct $K$-NN graphs [13], with $K=\{5,10,15\}$, and report the best performance. As features, we use the trained sequence encoder, which achieves better performance, used also in our method. In addition, we adopt ESM-1b as the residue encoder on GB1 dataset and adopt One-Hot encoding on AAV and Fluorescence datasets to speed up the training process. All experiments are conducted on two NVIDIA GeForce RTX 3090 GPUs and two NVIDIA RTX A6000 GPUs, and we report the mean performance of three runs with different random seeds. We present more details at https://github.com/zhiqiangzhongddu/EvolMPNN.
5.1 Effectiveness
EvolMPNN outperforms all baselines on 9 splits. Table 2 summarises performance comparison on AAV, GB1 and Fluorescence datasets. EvolMPNN achieves new state-of-the-art performance on most splits of three datasets, with up to $6.7\%$ improvements to baseline methods. This result vindicates the effectiveness of our proposed design to capture evolution information for homologous protein property prediction.
Table 2: Quality in terms of Spearman’s $\rho$ correlation with target value. NA indicates a non-applicable setting. * Used as a feature extractor with pre-trained weights frozen. $\dagger$ Results reported in [8, 44]. Top-2 performances of each split are marked as bold and underline.
Category Model Dataset
AAV GB1 Fluorescence
BLOSUM62 NA NA NA 0.128 0.005 -0.127 NA NA NA
DDE 0.649 0.636 0.158 0.445 0.816 0.306 0.690 0.638 0.159
Moran 0.437 0.398 0.069 0.069 0.589 0.193 0.445 0.400 0.046
Protein Seq. Encoder LSTM 0.125 0.608 0.308 -0.002 -0.002 -0.007 0.256 0.494 0.207
Transformer 0.681 0.748 0.304 0.271 0.877 0.474 0.250 0.643 0.161
CNN 0.746 0.730 0.406 0.502 0.857 0.515 0.805 0.682 0.249
ResNet 0.739 0.733 0.223 0.133 0.542 0.396 0.594 0.636 0.243
Pre-trained PLM ProtBert 0.794 0.719 0.322 0.634 0.866 0.308 0.451 0.679 0.201
ProtBert* 0.209 0.507 0.277 0.123 0.619 0.164 0.403 0.339 0.161
ESM-1b 0.821 0.735 0.385 0.704 0.878 0.386 0.804 0.679 0.221
ESM-1b* 0.454 0.573 0.241 0.337 0.605 0.178 0.528 0.430 0.091
ESM-1v 0.826 0.741 0.394 0.721 0.884 0.390 0.804 0.682 0.251
ESM-1v* 0.533 0.580 0.171 0.359 0.632 0.180 0.562 0.563 0.070
ESM-2 0.824 0.734 0.390 0.712 0.874 0.372 0.791 0.668 0.201
ESM-2* 0.475 0.581 0.199 0.422 0.632 0.189 0.501 0.511 0.084
GAT 0.821 0.741 0.369 0.757 0.873 0.508 0.768 0.667 0.208
GraphTransf. 0.827 0.749 0.389 0.753 0.876 0.548 0.780 0.678 0.231
NodeFormer 0.827 0.741 0.393 0.757 0.877 0.543 0.794 0.677 0.213
Ours EvolMPNN 0.835 0.757 0.433 0.768 0.889 0.584 0.809 0.684 0.262
Manual construction of homology graphs proves to be less effective. Notably, GNN-based methods that utilise manually constructed graph structure do not enter top-2 on 8 of 9 splits and two Transformer structure models, i.e., GraphTransformer and NodeFormer, often outperform such methods. It can be understood since homology graph construction is a challenging biomedical task [29], the simple $K$-NN graph construction is not an effective solution.
Large-scale PLM models are dominated by simple models. Surprisingly, we find that smaller models, such as CNN and ResNet, can outperform large ESM variants pre-trained on million- and billion-scale sequences. For instance, ESM-1v has about 650 million parameters and is pre-trained on around 138 million UniRef90 sequences [25]. Yet, CNN outperforms ESM-1v on three splits of Fluorescence dataset. This indicates the necessity of designs targeting specifically the crucial homologous protein engineering task.
Model Split
2-vs-R. 3-vs-R. L.-vs-H.
State-of-the-art 0.757 0.878 0.548
EvolMPNN ($\Phi^{\textsc{OH}}$) 0.766 0.877 0.553
EvolGNN ($\Phi^{\textsc{OH}}$) 0.764 0.866 0.536
EvolFormer ($\Phi^{\textsc{OH}}$) 0.764 0.868 0.537
EvolMPNN ($\Phi^{\textsc{PLM}}$) 0.768 0.881 0.584
EvolGNN ($\Phi^{\textsc{PLM}}$) 0.767 0.879 0.581
EvolFormer ($\Phi^{\textsc{PLM}}$) 0.766 0.879 0.575
5.2 Analysis of Performance
Figure 4: EvolMPNN performance on AAV’s Low-vs-High (a) and 2-vs-Rest (b) splits, with different hyper-parameters.
Influence of hyper-parameter settings on EvolMPNN. We present in Figure 4 a group of experiments to study the influence of some hyper-parameters on EvolMPNN, including the number of attention heads, embedding dimension and the number of layers of residue encoder and evolution encoder.
6 Conclusion and Future Work
We propose EvolMPNN that integrates both protein sequence information and evolution information by means of residues to predict the mutational effect on homologous proteins. Empirical and theoretical studies show that EvolMPNN and its extended variants (EvolGNN and EvolFormer) achieve outstanding performance on several benchmark datasets while retaining reasonable computation complexity. In future work, we intend to incorporate 3D protein structure information towards general-purpose homologous protein models. In addition, it would be interesting to experiment with different approaches to selecting anchor sets, for example, using central nodes-based selections.
Acknowledgments
Appendix 0.A Theoretical Analysis